OpenStudy (anonymous):

multiply 4xy/3y^4 * 9y/2x

OpenStudy (fallenangelorchid):

I'd really like to say I knew... But I don't xP Sorry @sleepyhead314

OpenStudy (holly00d1248):

Do you know @Legends

OpenStudy (anonymous):

thank you guys I appreciate the help =)

OpenStudy (anonymous):

Is this it? \(\dfrac{4xy}{3y^4} \times \dfrac{9y}{2x}\)

OpenStudy (anonymous):

If it is, we simplify it accordingly: \(\dfrac{4xy}{3y^4} \times \dfrac{9y}{2x} \rightarrow \dfrac{4xy \times 9y}{3y^4 \times 2x}\)

OpenStudy (anonymous):

yes =) how do you make it look like that on here? =o

OpenStudy (anonymous):

It's called LaTeX, just click the 'Equation' button below the comment box..it's a blue button.

OpenStudy (anonymous):

oh duh sorry I should have known that

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

Can you multiply 4xy * 9y? And 3y^4 * 2x?

OpenStudy (anonymous):

ohhh and yes =) .. I think lol 36xy^2 and 6xy^4 ?

OpenStudy (anonymous):

You got it! :D

OpenStudy (anonymous):

That gives us: \(\dfrac{36xy^2}{6xy^4}\) Now we divide the whole numbers and subtract the exponents: \(\dfrac{36}{6}x^{1-1}y^{2-4}\)

OpenStudy (anonymous):

Tell me what you get.

OpenStudy (anonymous):

6y^2 ? =)

OpenStudy (anonymous):

Yep! But the exponent is actually negative. \(6y^{-2}\)

OpenStudy (anonymous):

oh because the exponent placements in equation above and below the line?